Need help finding missing number
soverozou e %20B S % 20and % 20IS % 209628student %29%202015.xlsx File Edt Insert Format Heip Year 1 Year 2 Year 3 185,124 Year 4 Current assets Noncurrent assets 113,624 198,014 162,011 151,021 7 Total Assets Current liabilities Noncurrent liabilities 56,124 40.220 173,295 80,222 60,100 133,455 170,000 170,000 o Common stock Retained earnings 214,155 13,785 (3,644) Total liabilities and owners equity 524,600 388,456 1,452 13 4 Net Working Capital 150,222 140,000 165,000 15 16 In Handout
Answer #1

year 1

year 2

year 3

year 4

Current Assets

$   1,13,624.00

$ 1,90,442.00

$ 1,85,124.00

$ 2,45,222.00

Noncurrent assets

$   4,10,976.00

$ 1,98,014.00

$ 1,62,011.00

$ 1,51,021.00

Total Assets

$   5,24,600.00

$ 3,88,456.00

$ 3,47,135.00

$ 3,96,243.00

Current Liabilities

$       56,124.00

$      40,220.00

$ 1,15,583.00

$     80,222.00

Noncurrent liabilities

$   2,40,536.00

$ 1,73,295.00

$     60,100.00

$ 1,33,455.00

Common stock

$   2,14,155.00

$ 1,78,585.00

$ 1,70,000.00

$ 1,70,000.00

Retained earnings

$       13,785.00

$      (3,644.00)

$       1,452.00

$     12,566.00

Total liabilities and owner's equity

$   5,24,600.00

$ 3,88,456.00

$ 3,47,135.00

$ 3,96,243.00

Net Working Capital

$       57,500.00

$ 1,50,222.00

$ 1,40,000.00

$ 1,65,000.00

Working

year 1

year 2

year 3

year 4

Current Assets

113624

=388456-198014

185124

=165000+80222

Noncurrent assets

=524600-113624

198014

162011

151021

Total Assets

524600

388456

=185124+162011

=245222+151021

Current Liabilities

56124

40220

=347135-60100-170000-1452

80222

Noncurrent liabilities

=524600-56124-214155-13785

173295

60100

133455

Common stock

214155

=388456-40220-173295+3644

170000

170000

Retained earnings

13785

-3644

1452

=396243-80222-170000-133455

Total liabilities and owner's equity

524600

388456

347135

396243

Net Working Capital

=113624-56124

150222

140000

165000